How many liters of water are required to dissolve 1.00 g of silver chromate? Express your answer in liters to three significant figures. View Available Hint(s) nothing
Answer #1

Solution:

The solubility of silver chromate is 1.1 x10-12 M

Solubility is given in molarity (moles/litres). So to solve the problem we need to convert moles/litres to grams/litres. the molecular formula for silver chromate is Ag2CrO4, which has a molar mass of 332g/mole. so set up problem
( 1.1 x10-12 moles/1L) x (332g/1 mole) = (3.65×10^−10 g/L) (1g) x (L / 3.65x10^-10) = 2.74 x 109 L

therefore 2.74 x 109 L of water is required to dissolve 1.00 g of silver chromate.